OpenStudy (jonathanvaldes98):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = 5/3 y-intercept = (0, 2) slope = -5 y-intercept = (0, -6) slope = -3/5 y-intercept = (0, 2) slope = 5/3 y-intercept = (0, -6)

OpenStudy (anonymous):

Slope-intercept form is y = mx + b m is the slope {number in front of x} b is the y-intercept {number without a variable} 5x - 3y = -6 So what you do is subtract (5x from both sides) 3y = -5x+-6 Then you divide both sides (by 3) & You get slope = 5/3 y-intercept = (0, -6)

OpenStudy (jonathanvaldes98):

thankssssss alot ^ _ ^